

********************************************************************
UARS WEEKLY STATUS REPORT - 26 Dec 1997
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar observations on every
orbit and stellar calibration observations when possible.
***********
SUSIM
***********
Nothing received.
***********
PEM
***********
The PEM zenith spectrometers were active from December 4 through
December 7. PEM produced energy deposition profiles from electron and
proton precipitation from spectrometer data.
PEM is awaiting evaluation by GSFC of its proposed activation procedure.
The minimal configuration for PEM AUX activation is the zenith VMAG, MEPS1,
MEPS2, and HEPS1 units, only. This configuration is driven by the capabilities
of the spacecraft OBC and may increase if OBC program area increases. PEM
activation is proposed to be OBC driven when there exists plenty of power to
operate PEM. It is anticipated that the OBC will activate PEM on the daylight
portions of every orbit.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 12 December 1997 (Orbit 34178, GMT Day 346) through 19
December 1997 (Orbit 34282, GMT Day 353).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 347/0058 34178 AXIS on
HALOE 348/2326 34207 Off
PEM 350/2340 34237 Off
The beta angle reached a minimum angle of 0.0 degrees on 17 December 1997 (DOY
351). The beta angle is now increasing toward a maximum angle of 34.0 degrees
on 01 January 1998 (DOY 001).
A Reverse-to-Forward Yaw Around maneuver (number 64) was successfully executed
in Orbit 34246 at 14:41:26 UTC on 17 December 1997. The SSPP Instruments
(ACRIM, SOLSTICE, and SUSIM) were enabled. Due to the SA position at 269
degrees, the effective night length of the spacecraft after the maneuver was 48
minutes. Post maneuver analysis of data indicate that during the Yaw Around,
the end-of-night load bus voltage was 24.8 volts and the average depths of
discharge (DODs) on Batteries 2 and 3 were approximately 23.0 and 22.0 percent
respectively.
Two additional SSPP Solar Offset experiments were performed this week - UTC
Days 348 and 349. These experiments were identical to the normal solar offset
that is performed on a weekly basis.
On-Board Computer (OBC) system table definitions for tables 42 & 43 (star
tracker alignment & update filter status flags) were modified on Monday, 15
December 1997 (orbit 34220, 349/20:17:52 UTC). These changes were required to
allow clearing of update filter status telemetry flags which are sometimes set
during the yaw maneuvers (this functionality was inadvertently eliminated from
the update filter software included in flight software version 62). No code
changes were required, but due to updated system table definitions, the
on-orbit flight software is now designated version 62, revision C (V62C). An
operational database change was also required to support this revision (see
Ground System Operation).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 25.1 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.3 degrees C. Additional battery performance data for each battery on
13 December 1997 and 19 December 1997 are provided in Appendix D.
The clock error ranged from -10.0 to +9.7 msec during this report period.
Clock rate adjustments are listed in Appendix C.
On the 15 December 1997, a Rgyro glitchS occurred on the UARS spacecraft at
approximately 349/00:18 UTC. This resulted in an abrupt spacecraft pitch
movement of about 200 arc seconds error which slowly recovered to normal by
349/04:00 UTC. It took approximately four hours of star tracker and fine sun
sensor readings to bring the 200 arc second reference error back to normal and
allow the OBC reported attitude errors to again give accurate spacecraft
attitude telemetry. The last reported gyro glitch took place on 23 January
1996, and was discussed in UARS Anomaly Investigation Report AIR 96-010.
Further investigation into this gyro glitch continues.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 09
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT 01
% HALOE 01
% HRDI 03
% PEM 02
% SUSIM 01
% WINDII 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT 01
TRANSPONDER FREQUENCIES:
Transponder A: 97/343 - 2287.497007
Transponder B: 97/344 - 2287.499121
GROUND SYSTEM OPERATION
One (1) operational support problem(s) occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 97/352 show the observatory data loss to be 20 hours, 41
minutes, 12.916 seconds (no change). This is a 0.0378 percent data loss which
equals a 99.9622 percent data capture for the mission.
The FOT is still working out the details of participating in the IP return-link
telemetry test that the network is conducting.
Operational Database version 65 was released for use on Monday, 15 December
1997 (orbit 34220, 349/20:09:00 UTC). This release was scheduled to
correspond with changes to on-orbit system table definitions for tables 42 &
43 (star tracker alignment & update filter status flags) made during the same
contact (see Spacecraft Operation).
Flight Dynamics Facility (FDF) delivered a new Star Catalog to the FOT on 18
December. The new catalog was required to fix the invalid precession rate for
star 48000001. After the FOT verifies the integrity of the new catalog, it
will be configured for operational use on the Command Management System (CMS).
No testing of the new CMS software release (R7B1V2) occurred this week.
UARS Test and Training Simulator (UTTS) testing of Q-channel data flow was
conducted on 17 and 18 December 1997. Only a few OBC dumps were successfully
collected. On 17 December, the Q-channel driver did not indicate any data
flow. NASCOM and DOC saw clock data but no data across the Q-channel line.
The driver was switched and the first OBC data dump was received. However,
following attempts to dump the OBC failed. On 18 December, after
reconfiguring the CDH module the first dump of the OBC was good. However,
subsequent dumps were very sporadic and with many errors in the dump
comparisons. ATSC has been informed of the problem.
Operational processing of level-0 telemetry analysis files on the UMPGAL
computer (new MPG Alpha machine) was started on Monday, 15 December 1997. SSAI
corrected the problem with user account quotas on the UMPGAL computer, allowing
the successful processing of level-0 files on this machine. However, network
problems continued with transfer of the level-0 files from CDHF to UMPGAL,
preventing fully successful processing of these files. On Thursday, 18
December 1997, the FOT modified level-0 processing jobs on CDHF and UMPGAL to
perform verification checks on file transfer completion, and to perform
redundant transfers if necessary. Since then the level-0 processing jobs have
run successfully on UMPGAL in parallel with identical processing on CDHF.
OTHER SIGNIFICANT EVENTS
FOT Memo 97-051, RUARS Flight Software Version 62 Update Filter Changes,S was
released in response to numerous requests for this information from Code 512.
Monitoring of the magnetic field model performance continues.
FOT continues to improve a Power Profile for UARS based on the current status
of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
MEETINGS
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 16 January 1998 at
016/13:25:14 UTC.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust will
probably be required in mid-1998.
FACILITIES
no change
PERSONNEL CHANGES
no change
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 19 December 1997
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
34178 347/0058 97-622 PEM/AXIS Power ON
34187 347/1524 97-623 HRDI Heater delay to 150 amp. mins.
34190 347/2001 97-624 HRDI Heater delay to 50 amp. mins.
34203 348/1706 97-625 SUSIM Sec. Detector Off
34207 348/2326 97-626 HALOE Power OFF
34215 349/1214 97-627 HRDI Heater delay to 200 amp. mins.
34220 349/2009 97-628 OBC New system table 42 & 43 defs.
34235 350/2013 97-629 WINDII Heater delay to 350 amp. mins.
34237 350/2340 97-630 PEM Power OFF
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 19 December 1997
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRIFESURVTMP Y-HI HRDI IFE SURVIVAL TEMP 34179,83-4,98-99,214
HRRADTEMP Y-HI HRDI RADIOMETER TEMP 34179
WIEUTEMP Y-LO WINDII ELECTRONICS UNIT TEMP 34183-34282
OBSSOC2 Y-LO Battery 2 Current SOC 34185-219
OBSVLBEON Y-LO Load Bus Volts at EON 34189-206
OBSBATVEON3 Y-LO Battery 3 voltage at EON 34188-206
OBSBATVEON2 Y-LO Battery 2 voltage at EON 34190-206
QTSISAIF Y-LO SSPP Baseplate by SISA I/F 34192,258,61-65,73-82
SOSTMPRNTRN Y-LO Temperature Entrance 34273
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
(none)
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 19 December 1997
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
no changes remained at 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 19 December 1997
13 Dec 1997 - GMT Day 347 Beta = 18.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.62 +0.0/ +0.0 18.7 0.000 0.0
2 +33.6/ -22.4 +6.44/+5.15 +32.4/-15.2 25.1 1.024 22.9
3 +112.0/ +0.0 +4.83/+3.87 +31.2/-15.6 25.1 1.022 22.2
19 Dec 1997 - GMT Day 353 Beta = 4.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.62 +0.0/ +0.0 18.7 0.000 0.0
2 +44.8/ -22.4 +5.47/+4.51 +31.2/-13.6 25.1 1.023 21.8
3 +112.0/ +0.0 +4.51/+2.92 +29.6/-14.4 25.1 1.022 21.1
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
_______________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 19 December 1997
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
97-124 FOT 34210 349/0430 14170 Generic Late
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
SEVNTFW51